OpenStudy (anonymous):

Identify the curve by transforming the given polar equation to rectangular coordinates.

OpenStudy (anonymous):

letter (d)

OpenStudy (rational):

you only need to use below : \(x = r \cos \theta\) \(y = r \sin \theta \)

OpenStudy (rational):

and notice that \(x^2 + y^2 = r^2\)

OpenStudy (rational):

for #a : \(r = 2\) square both sides and replace r^2 by x^2+y^2, you're done.

OpenStudy (anonymous):

still kinda unsure how to go about this problem :/

OpenStudy (anonymous):

wait I'm sorry, i only need help with letter (d) !

OpenStudy (rational):

oh you're doing letter d.. okay lol :)

OpenStudy (anonymous):

hehe yes! thats the one :)

OpenStudy (rational):

cross multiply

OpenStudy (rational):

\[r = \dfrac{6}{3 \cos \theta + 2 \sin \theta}\] \[r(3 \cos \theta + 2 \sin \theta) =6\]

OpenStudy (anonymous):

mk...\[r(3\cos \theta+2\sin \theta)=3\]

OpenStudy (anonymous):

oops 6*

OpenStudy (rational):

distribute \(r\)

OpenStudy (anonymous):

let me do that real quick...!

OpenStudy (rational):

okie

OpenStudy (anonymous):

it looks a little weird :S ... let me write it

OpenStudy (anonymous):

\[r3\cos \theta+r2\sin \theta=6\]

OpenStudy (anonymous):

^?

OpenStudy (rational):

a*b = b*a so, r*3 = 3*r

OpenStudy (rational):

\[r = \dfrac{6}{3 \cos \theta + 2 \sin \theta}\] \[r(3 \cos \theta + 2 \sin \theta) =6\] \[ 3~r\cos \theta + 2~r \sin \theta =6\]

OpenStudy (anonymous):

thats all? :) @rational

OpenStudy (rational):

replace rcos and rsin by x and y

OpenStudy (rational):

\(r\) and \(\theta \) are polar variables, so you're not done until u have them in ur equation

OpenStudy (rational):

\[r = \dfrac{6}{3 \cos \theta + 2 \sin \theta}\] \[r(3 \cos \theta + 2 \sin \theta) =6\] \[ 3~r\cos \theta + 2~r \sin \theta =6\] \[ 3x+ 2y=6\]

OpenStudy (rational):

Now you're done. so basically the given polar curve is just a straight line !

OpenStudy (anonymous):

you were so helpful! thank you :D

OpenStudy (rational):

np :)